OpenStudy (anonymous):

A cone with a radius of 3" has a total area of 24 sq. in. Find the volume of the cone.

OpenStudy (adamaero):

need the volume equation for a cone first

OpenStudy (anonymous):

I did that. V=1/3pi(r^2)(h). But since you need the height of the altitude you need to find it by putting your equation for total area in, but you need to find you slant height first to find your altitude height.

OpenStudy (anonymous):

Then you use the PythagoreanTheorem

OpenStudy (adamaero):

um sure, or rearrange cone area = π*r(r+h^2+r^2) >> for h

OpenStudy (adamaero):

24/(3π) = 3 + h^2 + 9 8/π-12 = h^2 h = sqrt(8/π-12

OpenStudy (anonymous):

I found the answer. It is 12pi.